# A Multivariate Unimodality Test Harnessing the Dip Statistic of Mahalanobis Distances Over Random Projections

### v3

@@ Abstract
Unimodality, Dip, Random Projections, ToDo
@@ 1 Introduction
In parallel, another line of research has delved into the use of random projections as a strategy for capturing the essence of multi-dimensional distributions. Random projections, known for its efficacy in dimensionality reduction, have shown significant potential for learning mixtures of Gaussians (Dasgupta, 1999). Additionally, the Diaconis-Freedman effect elucidates the behavior of random projections of probability distributions in the high-dimensional space (Diaconis & Freedman, 1984). Specifically, for a given probability distribution $P$ in a $d$-dimensional space, when we consider a dimension $q$ much smaller than $d$, the majority of the $q$-dimensional projections of $P$ resemble scaled mixtures of spherically symmetric Gaussian distributions (Dümbgen et al., 2013). Consequently, random projections appear to be a potent tool for the analysis of unimodality as they facilitate the transformation of the problem into a seemingly simpler space.
In this paper, we introduce an algorithm intertwining random projections with the notion of $\alpha$-unimodality to devise a robust technique for multivariate unimodality testing. Central to our investigation is the extrapolation of one-dimensional unimodality principles to multi-dimensional domains via random projections and point-to-point distancing. Our method relies on the notion of $\alpha$-unimodality. We demonstrate that linear random projections preserve the $\alpha$-unimodality property in Mahalanobis distances from a reference point. Leveraging these one-dimensional Mahalanobis distances, we apply the dip test for unimodality detection. Employing various random projections, akin to Monte Carlo simulations, we assess the $\alpha$-unimodality of the original data distribution. In summary, the contribution of our work is two-fold. Firstly, we propose the first mathematically founded multivariate unimodality test, dubbed mud-pod [^1] Multivariate Unimodality Dip based on random Projections, Observers & Distances.. Secondly, we present the mp-means incremental clustering method which is a wrapper around k-means exploiting mudpod for unimodality assessment. Both theoretical findings and empirical validations underpin our methods, showcasing efficacy in unimodality assessments and clustering scenarios.
@@ 3.1 Preliminaries
In the following we introduce some definitions and additional notation that we will use throughout the paper. We use capital letters to denote random variables or matrices and boldface type to represent vectors.
@@ 3.1.1 $\alpha$-Unimodal Distrubutions
Without loss of generality, we assume that the mode of the unimodal distribution is at $\mathbf{0}$. A random d-vector $\mathbf{X}\in\mathbb{R}^{d}$ is said to have an $\alpha$-unimodal distrubution about $\mathbf{0}$ if, for every bounded, nonnegative, Borel measurable function $g$ on $\mathbb{R}^{d}$ the quantity $t^{\alpha}\operatorname{\mathbb{E}}[g(t\mathbf{X})]$ is nondecreasing in $t\in(0,\infty)$. In what follows, we use the notation $\mathbf{X}\sim\mathcal{P_{\alpha}}$ to represent a d-vector following an $\alpha$-unimodal distribution. It follows from the defintion that if $\mathbf{X}\sim\mathcal{P_{\alpha}}$ and $\alpha<\beta$, then $\mathbf{X}\sim\mathcal{P_{\beta}}$. An important result that provides an equivalent characterization for the set of $\alpha$-unimodal distrubutions is the Decomposition theorem: $\mathbf{X}\sim\mathcal{P_{\alpha}}$ iff $\mathbf{X}$ is distributed as $U^{\frac{1}{\alpha}}\mathbf{Z}$ where $U$ is uniform on $(0,1)$ and $\mathbf{Z}$ is independent of U (Olshen & Savage, 1970).
This theorem closely mirrors the intuition of one-dimensional unimodality, since Khintchine (1938) demonstrated that a real random variable $X$ has a unimodal distribution iff $X\sim UZ$, where $U$ is uniform on $[0,1]$ and $U$ and $Z$ are independent. It follows that a scalar $X\sim\mathcal{P_{\alpha}}$ iff $X^{\alpha}$ is unimodal as per the standard definition in $\mathbb{R}$. Next, we present and prove three pivotal properties of $\alpha$-unimodal distributions, i.e., the translation, norm, and projection properties.
Lemma 1 (Translation Property).
@@
    Let $t^{\alpha}\operatorname{\mathbb{E}}[g(t\mathbf{X}+t\mathbf{c})]=t^{\alpha}\operatorname{\mathbb{E}}[h(t\mathbf{X})]$, where $h(\mathbf{x})=g(\mathbf{x}+\mathbf{c})$. Note that $h$ is bounded, nonnegative, Borel measurable and that the first expression is nondecreasing iff the last expression is nondecreasing in $t$. $\blacksquare$
Lemma 2 (Norm Property).
Let $\mathbf{X}\sim\mathcal{P_{\alpha}}$, then $||\mathbf{X}||\sim\mathcal{P_{\alpha}}$.
@@
    $||\mathbf{X}||=\sqrt{(U^{\frac{1}{\alpha}}\mathbf{Z})^{\top}(U^{\frac{1}{\alpha}}\mathbf{Z})}=U^{\frac{1}{\alpha}}||\mathbf{Z}||$blacksquare
Lemma 3 (Projection Property).
@@
    A direct use of the Decomposition Theorem. $\blacksquare$
@@
    The proof arises from translation and projection lemmas. $\blacksquare$
The Mahalanobis distance possesses distinct properties: it is unitless, scale-invariant, and comprehensively considers the covariance structure across all dimensions. Traditionally, it has been employed in multivariate hypothesis testing. Notably, the Hotelling’s $T^{2}$ statistic (Hotelling, 1931), which generalizes the Student’s t-statistic, exemplifies its usage. In our study, the Mahalanobis distance is pivotal to our multivariate unimodality test.
@@ 3.1.2 Dip Test
The dip test serves as a tool for discerning multimodality within a unidimensional dataset. It gauges this by examining the maximum deviation between the empirical distribution function and the closest unimodal distribution function. The dip statistic for a distribution function $F$ is defined as: $\text{dip}(F)=\min_{G\in U}\rho(F,G)$, where $\rho(F,G)=\max_{t}|F(t)-G(t)|$ and $U$ represents the set of all possible unimodal distributions.
The dip test’s significance is highlighted by its ability to unveil the least among the most substantial deviations between the empirical cumulative distribution function (e.c.d.f.) $F$ of the univariate dataset and the c.d.f.s of the class of unimodal distributions. A salient attribute of the dip statistic is its convergence as the sample size burgeons, such that $\lim_{{n\to\infty}}\text{dip}(F_{n})=\text{dip}(F)$ (Hartigan & Hartigan, 1985). Moreover, the class of uniform distributions $U$ is acclaimed to be the most fitting for the null hypothesis, owing to its stochastically larger dip values compared to other unimodal distributions. The p-value for unimodality, derived via bootstrap samples, functions as a determinant for the dataset’s modality. A dataset with a p-value greater than $a$ indicates unimodality; otherwise, multimodality is suggested.
@@ 3.1.3 Random Projections
The Diaconis-Freedman effect is a valuable tool for unimodality analysis, simplifying the problem by likely transforming it into a Gaussian mixture model. When considering a probability distribution $P$ in a $d$-dimensional space with a lower-dimensional focus $q\ll d$, most $q$-dimensional projections of $P$ resemble scale mixtures of spherically symmetric Gaussian distributions. Additionally, linear random projections can approximate inter-point distances when projecting high-dimensional points into lower-dimensional spaces. This phenomenon is encapsulated in the celebrated Johnson & Lindenstraus lemma presented below (Johnson & Lindenstraus, 1984; Fernandez-Granda, 2016):
JL Lemma: Let $S:=\{x_{i}\}_{i=1}^{k}$ be a subset of $\mathbb{R}^{d}$. There exists a scaled random Gaussian matrix $\Pi\in\mathbb{R}^{d\times q}$ with $\Pi_{ij}\sim\mathcal{N}(0,1)$, where $m\geq\frac{8\log(k)}{\epsilon^{2}}$, ensuring for any $x_{i},x_{j}\in S$ with probability at least $\mathit{1/k}$:
$(1-\epsilon)\|x_{i}-x_{j}\|^{2}\leq\|\Pi x_{i}-\Pi x_{j}\|^{2}\leq(1+\epsilon)\|x_{i}-x_{j}\|^{2}$
We define $\mathcal{R}_{\mathbf{\Pi}}$ to be the set of all random projections satisfying the conditions of the Johnson–Lindenstrauss lemma. By employing a square root decomposition, we can demonstrate the applicability of the Johnson-Lindenstrauss lemma to the Mahalanobis distance (Bhattacharya et al., 2009), while mitigating the singularity issue inherent in inverting the covariance matrix in high dimensions (Lopes et al., 2011; Radhendushka Srivastava & Ruppert, 2016).
@@ 3.2 Connecting the Dots
where $\mathbf{\Sigma_{\Pi}}=\Pi\mathbf{\Sigma}\Pi^{T}$. It is important to note that given $\mathbf{X}\sim\mathcal{P_{\alpha}}$, both $\mathcal{D}_{\mathbf{S}}^{o}$ and $\Pi\circ\mathcal{D}_{\mathbf{S}}^{o}$ are $\alpha$-unimodal for any linear random projection $\Pi\in\mathcal{R}_{\mathbf{\Pi}}$, as established earlier. This yields the subsequent observation.
Proposition 1 (Randomisation Hypothesis).
Given the hypothesis that $\mathbf{X}\sim\mathcal{P_{\alpha}}$, $\mathcal{D}_{\mathbf{S}}^{o}$ is invariant under transformations in $\mathcal{R}_{\mathbf{\Pi}}$, ensuring that both $\mathcal{D}_{\mathbf{S}}^{o}$ and $\Pi\circ\mathcal{D}_{\mathbf{S}}^{o}$ exhibit an $\alpha$-unimodal distribution for any $\Pi\in\mathcal{R}_{\mathbf{\Pi}}$.
@@ 3.3 Multivariate Unimodality Testing
$H_{0}$ $:\mathbf{X}\sim\mathcal{P_{\alpha}}\quad\text{vs}\quad H_{1}:\mathbf{X}\not\sim\mathcal{P_{\alpha}}$
We define the pairing of a random projection $\Pi$ with an observer $\mathbf{o}$ as a random view. Given a set of $N$ random vectors $\mathcal{S}$ and a random view, we can obtain the corresponding set of Mahalanobis distances $\Pi\circ\mathcal{D}_{\mathbf{S}}^{o}$. Under the null hypothesis, recall that $\Pi\circ\mathcal{D}_{\mathbf{S}}^{o}=\{d_{i}\}_{i=1}^{N-1}$ is $\alpha$-unimodal, and the set $\{d_{i}^{\alpha}\}_{i=1}^{N-1}$ is unimodal, allowing the employment of the dip test. Let $T(\Pi\circ\mathcal{D}_{\mathbf{S}}^{o})$ denote the dip test p-value. If $a\in[0,1]$ is the significance level, the null hypothesis is rejected iff $T(\Pi\circ\mathcal{D}_{\mathbf{S}}^{o})\leq a$.
Utilizing the idea of random views, which, as previously discussed, preserve $\alpha$-unimodality, we can employ them as a foundation for Monte Carlo simulations. The rationale is that as more views reject the null hypothesis, our confidence about data multimodality increases. Let $\{\Pi_{i}\}_{i=1}^{M}\subset\mathcal{R}_{\mathbf{\Pi}}$ be a set of $M$ random projections. Leveraging Monte Carlo hypothesis testing theory (Lehmann & Romano, 2005, Chapter 11.2.2) and building on the fact that the dip statistic has a well defined c.d.f. (Hartigan & Hartigan, 1985), we explore the conditional c.d.f. of the dip test: $J_{N}(t)=P\left(T(\Pi\circ\mathcal{D}_{\mathbf{S}}^{o})\leq t\mid\mathcal{D}_{\mathbf{S}}^{o})\right)$. The aforementioned probability is measured over the joint distribution of random projections $\mathcal{R}_{\mathbf{\Pi}}$ and the distribution of observers $\mathcal{O}$. Let ${.}I\{.\}$ denote the indicator function, we define $\hat{J}_{n,M}(t)$ as the approximation of the true c.d.f $J_{N}(t)$ computed on the series of the random views:
@@
Hitherto, we have not delineated the methodology for selecting observers from the set of points obtained from a random projection. Several sampling strategies can exist, with the most intuitive being uniform random sampling. However, empirical results suggest that uniformly selecting observers based on a specific percentile of the distance distribution from the samples’ mean yields superior performance. The underlying rationale is that points situated farther away from the means possess a better capability to discern the topographical elevations formed by distinct clusters (Kalogeratos & Likas, 2012). Ultimately, the projection dimension is the minimal integer that satisfies the JL lemma for a specified $\epsilon$. For a given distribution family $\mathcal{P_{\alpha}}$, the complete mud-pod test is detailed in Algorithm 1.
@@
7:  return $p=\frac{1}{M}\sum_{i=1}^{M}I\{T(\Pi_{i}\circ\mathcal{D}_{\mathbf{S}}^{o})\leq a\}$
It is important to note that dip-dist (Kalogeratos & Likas, 2012) is a special case of mud-pod, omitting $\alpha$ exponent, operating in the original space using the Euclidean distance and considering all data points as observers.

### v4

@@ Abstract
multivariate unimodality test, alpha-unimodality, dip test, random projections, clustering, cluster number estimation
@@ 1 Introduction
In parallel, another line of research has delved into the use of random projections as a strategy for capturing the essence of multi-dimensional distributions. Random projections, known for its efficacy in dimensionality reduction, have shown significant potential for learning mixtures of Gaussians (Dasgupta, 1999). Additionally, the Diaconis-Freedman effect elucidates the behavior of random projections of probability distributions in the high-dimensional space (Diaconis & Freedman, 1984). Specifically, for a given probability distribution $P$ in a $d$-dimensional space, when we consider a dimension $q$ much smaller than $d$, the majority of the $q$-dimensional projections of $P$ resemble scaled mixtures of spherically symmetric Gaussian distributions (Dümbgen et al., 2013). Consequently, random projections appear to be a potent tool for the analysis of unimodality as they facilitate the transformation of the problem into a seemingly simpler space. However, not every random projection is conducive to our analysis; many projections can obfuscate distinct modes by distorting distances. Consequently, we limit our approach to a family of random projections that, with a certain probability, maintain pairwise distances.
In this work, we tackle the complexity of extending unimodality testing to higher dimensions. Echoing the principles of $\alpha$-unimodality, we introduce a novel algorithm for efficient multivariate unimodality testing. Our approach bridges the gap between the simplicity of one-dimensional unimodality confirmation and the intricacies of its higher-dimensional counterpart. Central to our investigation are the $\alpha$-unimodality preserving properties of point-to-point distancing and linear random projections. We demonstrate that linear random projections preserve the $\alpha$-unimodality property in Mahalanobis distances from a reference point. Leveraging these one-dimensional Mahalanobis distances, we apply the dip test for unimodality detection. Employing various random projections, akin to Monte Carlo simulations, we assess the $\alpha$-unimodality of the original data distribution. In summary, the contribution of our work is two-fold. Firstly, to the best of our knowledge, we propose the first mathematically founded multivariate unimodality test, dubbed mud-pod [^1] Multivariate Unimodality Dip based on random Projections, Observers & Distances.. Secondly, we present the mp-means incremental clustering method which is a wrapper around k-means exploiting mudpod for unimodality assessment. Both theoretical findings and empirical validations underpin our methods, showcasing efficacy in unimodality assessments and clustering scenarios.
@@ 3.1 Preliminaries
In the following, we present key notation and foundational background applicable throughout this paper. We use capital letters to denote random variables or matrices and boldface type to represent vectors.
@@ 3.1.1 $\alpha$-Unimodal Distrubutions
Without loss of generality, we assume that the mode of the unimodal distribution is at $\mathbf{0}$. A random d-vector $\mathbf{X}\in\mathbb{R}^{d}$ is said to have an $\alpha$-unimodal distrubution about $\mathbf{0}$ if, for every bounded, nonnegative, Borel measurable function $g$ on $\mathbb{R}^{d}$ the quantity $t^{\alpha}\mathbb{E}[g(t\mathbf{X})]$ is nondecreasing in $t\in(0,\infty)$. In what follows, we use the notation $\mathbf{X}\sim\mathcal{P_{\alpha}}$ to represent a d-vector following an $\alpha$-unimodal distribution. It follows from the defintion that if $\mathbf{X}\sim\mathcal{P_{\alpha}}$ and $\alpha<\beta$, then $\mathbf{X}\sim\mathcal{P_{\beta}}$. An important equivalent characterization for the set of $\alpha$-unimodal distrubutions is the Decomposition theorem: $\mathbf{X}\sim\mathcal{P_{\alpha}}$ iff $\mathbf{X}$ is distributed as $U^{\frac{1}{\alpha}}\mathbf{Z}$ where $U$ is uniform on $(0,1)$ and $\mathbf{Z}$ is independent of $U$ (Olshen & Savage, 1970).
This theorem closely mirrors the intuition of one-dimensional unimodality, since Khintchine (1938) demonstrated that a real random variable $X$ has a unimodal distribution iff $X\sim UZ$, where $U$ is uniform on $[0,1]$ and $U$ and $Z$ are independent. It follows that a scalar $X\sim\mathcal{P_{\alpha}}$ iff $X^{\alpha}$ is unimodal as per the standard definition in $\mathbb{R}$. Next, we present and prove three pivotal properties of $\alpha$-unimodality, i.e., the translation, norm, and projection properties.
Lemma 3.1 (Translation Property).
@@
Let $t^{\alpha}\mathbb{E}[g(t\mathbf{X}+t\mathbf{c})]=t^{\alpha}\mathbb{E}[h(t\mathbf{X})]$, where $h(\mathbf{x})=g(\mathbf{x}+\mathbf{c})$. Note that $h$ is bounded, nonnegative, Borel measurable and that the first expression is nondecreasing iff the last expression is nondecreasing in $t$. ∎
Lemma 3.2 (Norm Property).
If $\mathbf{X}\sim\mathcal{P_{\alpha}}$, then $||\mathbf{X}||\sim\mathcal{P_{\alpha}}$.
@@
$||\mathbf{X}||=\sqrt{(U^{\frac{1}{\alpha}}\mathbf{Z})^{\top}(U^{\frac{1}{\alpha}}\mathbf{Z})}=U^{\frac{1}{\alpha}}||\mathbf{Z}||$
Lemma 3.3 (Projection Property).
@@
A direct use of the Decomposition Theorem. ∎
@@
Proof stems from translation and projection lemmas. ∎
Mahalanobis distance possesses distinct properties: it is unitless, scale-invariant, and considers the covariance structure across all dimensions. Traditionally, it has been employed in multivariate hypothesis testing. Notably, the Hotelling’s $T^{2}$ statistic (Hotelling, 1931), which generalizes the Student’s t-statistic, exemplifies its usage. The Mahalanobis distance is pivotal to our multivariate unimodality test.
@@ 3.1.2 Dip Test
The dip test serves as a tool for discerning multimodality within a unidimensional dataset. It gauges this by examining the maximum deviation, i.e., the Kolmogorov-Smirnov statistic, between the empirical cumulative distribution function (e.c.d.f.), $F(t)$, and the nearest unimodal c.d.f., $G(t)$. The dip statistic for a distribution function $F$ is defined as: $\text{dip}(F)=\min_{G\in U}\rho(F,G)$, where $\rho(F,G)=\max_{t}|F(t)-G(t)|$ and $U$ represents the set of all possible unimodal distributions. The dip test’s significance is highlighted by its ability to unveil the least among the most substantial deviations between the empirical cumulative distribution function $F$ of the univariate dataset and the c.d.f.s of the class of unimodal distributions. A salient attribute of the dip statistic is its convergence as the sample size burgeons, such that $\lim_{{n\to\infty}}\text{dip}(F_{n})=\text{dip}(F)$ (Hartigan & Hartigan, 1985). Moreover, the class of uniform distributions $U$ is acclaimed to be the most fitting for the null hypothesis, owing to its stochastically larger dip values compared to other unimodal distributions. To calculate the dip value, the e.c.d.f. of the data is considered, and the unimodal piecewise linear function with the smallest maximum distance to the e.c.d.f. is determined. The p-value for unimodality, derived via bootstrap samples, functions as a determinant for the dataset’s modality. A dataset with a p-value greater than $a$ indicates unimodality; otherwise, multimodality is suggested.
@@ 3.1.3 Random Projections
The Diaconis-Freedman effect can be a valuable tool for unimodality analysis, simplifying the problem by likely transforming it into a Gaussian mixture model. When considering a probability distribution $P$ in a $d$-dimensional space, most $q$-dimensional projections of $P$ with $q\ll d$ resemble scale mixtures of spherically symmetric Gaussian distributions. Additionally, linear random projections can preserve distances when projecting high-dimensional points into lower-dimensional spaces. This phenomenon is encapsulated in the celebrated Johnson & Lindenstraus lemma presented below (Johnson & Lindenstraus, 1984; Fernandez-Granda, 2016):
JL Lemma: Let $S:=\{x_{i}\}_{i=1}^{k}$ be a subset of $\mathbb{R}^{d}$ and $\epsilon>0$. Then, let $\Pi\in\mathbb{R}^{d\times q}$, where $q\geq\mathit{8\log(k)/\epsilon^{2}}$, be a random matrix with i.i.d. entries $\Pi_{ij}\sim\mathcal{N}(0,\mathit{1/d)}$. With probability at least $\mathit{1/k}$, for any $x_{i},x_{j}\in S$, we have:
$(1-\epsilon)\|x_{i}-x_{j}\|^{2}\leq\|\Pi x_{i}-\Pi x_{j}\|^{2}\leq(1+\epsilon)\|x_{i}-x_{j}\|^{2}.$
We denote by $\mathcal{R}_{\mathbf{\Pi}}$ the set of matrices fulfilling the distance preservation criteria specified in the JL Lemma. According to the JL Lemma, if $\Pi$ is sampled with i.i.d. $\mathcal{N}(0,\frac{1}{d})$ entries, then $P(\Pi\in\mathcal{R}_{\Pi})\geq\frac{1}{k}$. By employing a square root decomposition, we can demonstrate the applicability of the JL lemma to the Mahalanobis distance (Bhattacharya et al., 2009), while mitigating the singularity issue inherent in inverting the covariance matrix in high dimensions (Lopes et al., 2011; Radhendushka Srivastava & Ruppert, 2016).
@@ 3.2 Connecting the Dots
where $\mathbf{\Sigma_{\Pi}}=\Pi\mathbf{\Sigma}\Pi^{T}$. It is important to note that since $\mathbf{X}\sim\mathcal{P_{\alpha}}$, the elements of both $\mathcal{D}_{\mathbf{S}}^{o}$ and $\Pi\circ\mathcal{D}_{\mathbf{S}}^{o}$ exhibit $\alpha$-unimodal distributions, as established earlier. This yields the subsequent observation.
Proposition 3.5 (Randomisation Hypothesis).
Given $\mathbf{X}\sim\mathcal{P}_{\alpha}$, the distributions of elements within $\mathcal{D}_{\mathbf{S}}^{o}$ and $\Pi\circ\mathcal{D}_{\mathbf{S}}^{o}$ retain $\alpha$-unimodality under any transformation $\Pi\in\mathcal{R}_{\mathbf{\Pi}}$.
@@ 3.3 Multivariate Unimodality Testing
$H_{0}$ $:\mathbf{X}\sim\mathcal{P_{\alpha}}\quad\text{vs.}\quad H_{1}:\mathbf{X}\not\sim\mathcal{P_{\alpha}}.$
We define the pairing of a random projection $\Pi$ with an observer $\mathbf{o}$ as a random view. We assume independence between the random projection $\Pi$ and the observer $\mathbf{o}$. Given a set of $N$ random vectors $\mathcal{S}$ and a random view, we can obtain the corresponding set of Mahalanobis distances $\Pi\circ\mathcal{D}_{\mathbf{S}}^{o}$. Under the null hypothesis, recall that $\Pi\circ\mathcal{D}_{\mathbf{S}}^{o}=\{d_{i}\}_{i=1}^{N-1}$ is $\alpha$-unimodal, and the set $\{d_{i}^{\alpha}\}_{i=1}^{N-1}$ is unimodal, allowing the employment of the dip test. Let $T(\Pi\circ\mathcal{D}_{\mathbf{S}}^{o})$ denote the dip test p-value. If $a\in[0,1]$ is the significance level, the null hypothesis is rejected iff $T(\Pi\circ\mathcal{D}_{\mathbf{S}}^{o})\leq a$.
Utilizing the idea of random views, which, as previously discussed, preserve $\alpha$-unimodality, we can employ them as a foundation for Monte Carlo simulations. The rationale is that as more views reject the null hypothesis, our confidence about data multimodality increases. Let $\{\Pi_{i}\}_{i=1}^{M}\subset\mathcal{R}_{\mathbf{\Pi}}$ be a set of $M$ random projections. Leveraging Monte Carlo hypothesis testing theory (Lehmann & Romano, 2005, Chapter 11.2.2) and building on the fact that the dip statistic has a well defined c.d.f. (Hartigan & Hartigan, 1985), we explore the conditional c.d.f. of the dip test: $J_{N}(t)=P\left(T(\Pi\circ\mathcal{D}_{\mathbf{S}}^{o})\leq t\mid\mathcal{S}_{\mathbf{X}})\right)$. The aforementioned probability is measured over the joint distribution of random projections $\mathcal{R}_{\mathbf{\Pi}}$ and the distribution of observers $\mathcal{O}$. Let ${.}I\{.\}$ denote the indicator function, we define $\hat{J}_{n,M}(t)$ as the approximation of the true c.d.f $J_{N}(t)$ computed on the series of the random views:
@@
Hitherto, we have not delineated the methodology for selecting observers from the set of points obtained from a random projection. Several sampling strategies can exist, with the most intuitive being uniform random sampling. However, empirical results suggest that uniformly selecting observers based on a specific percentile of the distance distribution from the samples’ mean yields superior performance. The underlying rationale is that points situated farther away from the means possess a better capability to discern the topographical elevations formed by distinct clusters (Kalogeratos & Likas, 2012). It is important to note that despite the dependency introduced between the observer $\mathbf{o}$ and the random projection $\Pi$ by the percentile strategy, our analysis remains valid thanks to the extension of the Glivenko-Cantelli theorem to strictly stationary sequences (Sharipov, 2011). Ultimately, the projection dimension is the minimal integer that satisfies the JL lemma for a specified $\epsilon$. Algorithm 1 details the complete mud-pod test.
@@
7:  return $p_{\text{value}}=\frac{1}{M}\sum_{i=1}^{M}I\{T(\Pi_{i}\circ\mathcal{D}_{\mathbf{S}}^{o})\leq a\}$